# HalluciBot: Is There No Such Thing as a Bad Question?

### v1

@@ 4 Metrics for Hallucination
Reliability
We employ Cronbach’s $\alpha$ to assess the internal consistency of our Monte Carlo Simulation, as a corpus-level statistic (Cronbach, 1951).
Scenario Accuracy Agreement Rel
@@
WAVG Extractive 85,892 92.1 91.0 69.0 97.4 87.2 85.5 84.3 75.3 99.8
Multiple Choice 81,697 76.3 76.8 47.4 91.6 71.8 75.2 71.3 61.9 92.5
Abstractive 202,248 55.9 53.9 32.9 67.3 51.2 81.5 80.0 69.1 98.8
Total 369,837 68.6 67.4 44.3 79.4 63.9 81.0 79.1 69.0 97.7
Table 3: Stage 1 Monte Carlo: Weighted average (WAVG) results are presented for each scenario and in total. The metrics are calculated using data from all examples within each scenario, regardless of the original dataset split. For a more detailed breakdown, refer to Table 15.
5 Experimental Setup
5.1 Dataset Coverage & Scenario Split
Our experiment includes 13 datasets (Table 11) divided into 3 scenarios: Extractive, Multiple Choice, and Abstractive. To evaluate the impact of context, we use SQuADv2 (Rajpurkar et al., 2016, 2018) to simulate Retrieval-Augmented Generation (Lewis et al., 2021; Guu et al., 2020). To assess the effect of multiple choice queries, we repurposed TruthfulQA (Lin et al., 2022) and SciQ (Johannes Welbl, 2017) for two experiments: one where the LLM agents select from the choices or context, and another where LLM agents generate outputs without context. Prompt templates for each output agent can be found in Table 8. We maintain the original train, validation, and test splits for each dataset across scenarios to prevent information leakage to HalluciBot.
5.2 Monte Carlo Parameters & Setup
Our Monte Carlo Simulation utilizes 2 LLM components: the Query Perturbator and Output Generator. We use gpt-3.5-turbo to sample perturbations and outputs based on the prompts provided in Table 8. All LLM agents share the same set of parameters, as described in Table 6.
Well-formedness
We used a syntactically-aware well-formedness scoring RoBERTa model (Kumar, 2021) trained on the Google Query Well-formedness Dataset (Faruqui and Das, 2018) to evaluate the grammatical correctness and completeness of 1,881,005 synthetically generated queries.
5.3 HalluciBot Parameters & Environment
We employed HuggingFace’s Trainer class with the Adam optimizer (Kingma and Ba, 2017) for training, reporting efficiency and training times in Table 9. All experiments were conducted on an AWS EC2 instance with a single GPU (Table 7).
Base Models
We fine-tune pretrained BERT (Devlin et al., 2018) and RoBERTa (Liu et al., 2019) models (Table 10). To address label imbalance, we employed a weighted loss where each class weight is assigned to its inverted frequency in the training set. For the multi-class predictor, we trained a cross-entropy and ordinal model.
Train Splits
The train, validation, and test splits follow the original divisions of the datasets. Specifically, there are 302,492 training, 44,491 validation, and 22,854 testing samples. The distribution of labels across these splits is summarized in Table 2, and fine-grained splits per set are in Table 16.
@@ 5.3 HalluciBot Parameters & Environment
+ Scenario 84.9 72.9 68.6 85.0 76.9 71.1 92.1 80.9 78.2 78.8 73.2 65.3
   + $\tau=0.146$ 82.5 72.5 69.2 85.0 79.5 75.3 79.0 73.4 71.8 92.0 86.8 79.3
Table 4: HalluciBot Binary Evaluation Statistics. We report the Accuracy, F1, Precision, and Recall for all data splits. Probability threshold $\tau$ is computed along the closed interval $[0,1]$ in increments of $0.001$ to maximize the F1 score. The best ablation per base model is underlined, while the overall best performing model is in bold.
@@
Model Train Val Test Train Val Test Train Val Test Train Val Test
BERT-base-cased 49.6 32.2 24.7 69.7 49.2 40.7 81.4 62.7 56.4 70.6 53.2 49.6
+ Scenario 54.1 38.7 31.3 72.2 54.8 46.1 82.8 67.6 59.3 76.4 60.6 54.3
+ Ordinal 58.7 45.3 38.6 70.0 54.5 48.3 79.0 64.1 59.1 75.9 61.4 55.8
RoBERTa-base 47.6 34.1 26.6 66.2 50.1 42.6 77.9 62.7 57.3 67.6 54.1 51.5
+ Scenario 52.2 41.5 34.4 69.2 57.0 48.4 79.8 68.6 59.5 73.5 62.6 56.1
+ Ordinal 47.8 39.4 37.1 56.7 48.7 46.6 67.0 60.0 57.8 64.1 56.1 54.4
RoBERTa-large
+ Scenario 61.6 47.6 38.8 77.5 62.6 53.1 85.8 73.0 63.8 79.5 65.4 58.8
+ Ordinal 60.8 48.0 40.7 73.6 59.0 52.2 81.9 67.5 62.3 76.7 63.4 57.4
Table 5: HalluciBot Multi-class Evaluation Statistics. Considering the challenge of approximating a random variable and the potential presence of noise in our empirical estimate, we provide accuracy measurements for Top 1, Top 2, and Top 3 predictions. Additionally, we report the accuracy within a margin of ±$\pm 1$ of the predicted class.
@@ 6 Analysis & Discussion
6.1 Aggregated Monte Carlo Results
The aggregated Monte Carlo results demonstrate the relative performance of different question-answering scenarios. On average, the results (Table 3) indicate that Extractive outperforms Multiple Choice, which, in turn, outperforms Abstractive when subjected to perturbations. This trend suggests that the performance of gpt-3.5-turbo is influenced by the presence of additional content. Abstractive tasks show the greatest variation in agent response under perturbations, highlighting the effectiveness of added context in mitigating hallucinations (Figures 3 & 4). A full breakdown of our results is provided in Table 15.
Extractive QA
With context, gpt-3.5-turbo performs well on the SQuADv2 dataset. The mode accuracy (91.0%) and agreement (75.3%) of the agents are high. Even under radical perturbations, having unaltered context provides robustness to the agent’s capacity to answer correctly.
Multiple Choice QA
Access to answer choices mitigates hallucinations across perturbations. The ensemble accuracy is slightly higher than the baseline (+0.5%), showcasing that multiple agents can (slightly) improve accuracy rates.
Abstractive QA
When no additional context is provided, gpt-3.5-turbo achieves a mode accuracy of 53.9% under perturbations. Interestingly, there is a significant dispersion of hallucination rates compared to other scenarios (Figure 3). Moreover, there is significant variation in results among datasets. For instance, SQuADv2 shows a -59% decrease in baseline accuracy against its Extractive counterpart. In contrast, SciQ benefits in this setting, leading to a +9.4% increase in mode accuracy, as the likelihood of generating a match increases.
Coherence of Perturbations
6.2 Can HalluciBot detect Hallucination?
Binary Results
The best HalluciBot achieves a validation accuracy of 73.0%, with a macro F1-score of 76.9% (RoBERTa-base + Scenario). With threshold tuning, improved recall and precision rates yield a F1-score of 79.6%. A full breakdown of our results is provided in Table 4.
Multi-class Results
HalluciBot achieves a validation accuracy of 47.6%, with a Top 3 accuracy of 73.0% for the RoBERTa-large + Scenario model.
Does Knowing the Scenario Matter?
Including the scenario in HalluciBot’s prompt has proven beneficial in distinguishing similar queries with varying contexts. For binary cases, incorporating the scenario leads to a notable increase of +15.5% in F1. For multi-class validation, the inclusion of scenario results in a +21.7% increase in Top 1 Accuracy and a +9.4% increase in Top 3 Accuracy.
Test Distribution Shift
The testing set poses three challenges: (1) the absence of Extractive samples; (2) an over-representation of Multiple Choice queries (94.4%); and (3) a distribution of multi-class labels that diverges with significantly fewer hallucinating-inducing queries compared to the training and validation sets. Consequently, we rely on the validation results instead.

### v3

@@ 2 Related Work
3 Methodology Overview
What is Hallucination? In general terms, hallucination refers to a false perception of patterns or objects resulting from one’s senses. With regards to LLMs, a myriad of studies bifurcate into (1) factuality hallucinations that refer to outputs which directly contradict or fabricate the ground truth while (2) faithfulness hallucinations define outputs that misunderstand the context or intent of the query (Huang et al. 2023; Ji et al. 2023). In this study, we introduce truthful hallucination as the motivation on why we are perturbing the original query. Truthful hallucination is defined as an LLM’s inability to answer semantically similar but lexically different perturbations of a query. The motivation for truthful hallucination stems from the analysis that neural networks display an intrinsic propensity to memorize training data (Carlini et al. 2021) - in this case, memorizing the query and output. Given the risk of over-training LLMs, their opaque training data, and propensity to memorize - generating multiple outputs from the same query or analyzing a single output from a single query do not help measure truthful hallucination.
What is the Motivation for HalluciBot? HalluciBot focuses on distilling LLM behavior into a speedy encoder that can predict hallucination before generation. Foremost, this is in contrast to prior work that uses multiple generations during a user’s session to provide self-consistency (Manakul, Liusie, and Gales 2023). Next, our proposal differs from entropy based, log-prob based, or model based estimation techniques (Huang et al. 2023) that rely on the LLM’s uncertainty to predict hallucinations - these methods focus on the model’s bias while we focus on empirical estimates. Moreover, our approach consists of a Multi-Agent simulation which stands in stark contrast to the majority of current experiments that have focused on leveraging a single LLM agent to generate outputs from a single query (Cole et al. 2023; Kadavath et al. 2022; Snyder, Moisescu, and Zafar 2023). The training procedure for HalluciBot consists of perturbing each query $n=5$ times, employing $n+1=6$ independent LLM agents, sampling an output from each query, conducting a Monte Carlo simulation on 2,219,022 sampled outputs, and training an encoder classifier.
3.1 Multi-Agent Monte Carlo Simulation
What is the Purpose of a Monte Carlo Simulation? As evidenced by multiple studies and Table 3, hallucination is the outcome of multiple confounding variables - thus, it is highly unlikely that a tractable closed-form solution will be able to model hallucinations. Thus, we employ a Monte Carlo simulation as a means to derive empirical estimations of hallucination rates in LLMs, since this method is frequently leveraged to map probability in the presence of random variable inference (Swaminathan 2021). Thus, we estimate the probability density that a query induces hallucination.
What is a Query Perturbator? Via perturbations, we induce diversity to disentangle the generation process from any potential training bias (Alzahrani et al. 2024; Carlini et al. 2021). The Query Perturbator is a gpt-3.5-turbo LLM agent that generates $n=5$ perturbations to the original query $q_{0}$ while retaining the same semantic meaning. In effect, the generation process can be summarized as returning a set of $\mathcal{Q}=\left\{q_{0},q_{1},\ldots,q_{n}\right\}$ query perturbations of size $n+1$. The Query Perturbator’s singular purpose is to: Rewrite the query in {$n$} radically different ways. One prompt call is sufficient to discourage duplicates. Temperature is set to $1.0$ to prioritize creativity and lexical diversity. Our analysis in Table 3 shows that introducing perturbations, rather than sampling $n+1$ outputs for query $q_{0}$, results in a $13$ point spread between the lower and upper bound accuracy, a $12.5$ point decrease in Fleiss’s $\kappa$ for agreement, while the modal accuracy remains largely unchanged. This suggests that perturbations inject variability into our Monte Carlo simulation, which is critical for observing diverse outputs and hallucinations. This corroborates the observation by Alzahrani et al. (2024) that LLMs are highly sensitive to even minor details.
What is an Output Generator? For the perturbed set $\mathcal{Q}$ for a sample $q_{0}$, the Output Generator consists of $|\mathcal{Q}|=n+1$ six independent gpt-3.5-turbo LLM agents to generate outputs $a_{i}\in\mathcal{A}$ for each variation $q_{i}\in\mathcal{Q}$. The LLM agent will receive (1) for Extractive queries, a prompt with the query $q_{i}$, alongside context $c_{i}$, (2) for Multiple-Choice queries, candidate choices $k_{i}\in\mathcal{K}$, and (3) for Abstractive queries, no additional context. Table 8 outline’s each experiment’s prompt procedure. Temperature for all experiments is set to $1.0$ to stress-test and encourage diversity.
How Do We Measure Accuracy? Accuracy serves as the measure of correctness, comparing the generated output $a_{i}$ to the ground truth $y$, using partial, case-insensitive matching with the TheFuzz library. For Multiple Choice queries, the choice label is also considered. If there is no match between the output $a_{i}$ and the ground truth $y$, we assign $\mathbb{I}\bigl{[}a_{i}\not=y\bigr{]}\mapsto 1$; otherwise, $\mathbb{I}\bigl{[}a_{i}=y\bigr{]}\mapsto 0$. The results are compared to the baseline (original query $q_{0}$, output $a_{0}$), the mode (most common $a_{i}$), the lower bound (all correct), and the upper bound (at least one $a_{i}$ correct).
How Do We Measure Agreement? Accuracy alone is insufficient for evaluating the agreement among multiple agents. To assess agreement, we report statistical measures for our Monte Carlo experiments including Item Difficulty ($\mathbf{\mu_{D}}$) (Lord 1952), Fleiss’s Generalized $\mathbf{\kappa}$ (Cohen 1960; Fleiss 1971), Mean Certainty / Entropy ($\mathbf{H_{\eta}}$) (Shannon 1948; Wilcox 1973), and Gibbs’ $\mathbf{M_{2}}$ Index (Gibbs and Poston 1975). These metrics help evaluate agreement levels amongst independent agents. For instance, high agreement on an incorrect answer indicates a misconception, while low agreement could suggest confusion or a poorly formulated query. To address this limitation in HalluciBot, we introduce a dual cross-entropy loss based on hallucination rates and consensus to improve the model’s ability to distinguish good queries from bad queries.
Scenario Accuracy Agreement
@@ 3.1 Multi-Agent Monte Carlo Simulation
SINGLE Extractive 85,734 89.8 90.3 83.6 94.6 89.8 91.4 92.0 90.4
Multiple Choice 80,813 74.0 75.8 58.1 88.0 73.8 90.3 83.7 75.5
Abstractive 200,693 56.2 56.7 44.2 67.4 56.1 93.2 89.8 80.2
Total 367,240 68.0 68.7 56.4 78.3 67.9 94.4 90.2 81.5
MULTI Extractive 85,892 92.1 91.0 69.0 97.4 87.2 85.5 84.3 75.3
Multiple Choice 81,697 76.3 76.8 47.4 91.6 71.8 75.2 71.3 61.9
Abstractive 202,248 55.9 53.9 32.9 67.3 51.2 81.5 80.0 69.1
Total 369,837 68.6 67.4 44.3 79.4 63.9 81.0 79.1 69.0
Table 3: Comparing Single Query, Multiple Outputs (SINGLE) vs. Single Query, Multiple Perturbations, Single Output (MULTI) Monte Carlo Experiments (§5). The reported metrics (§3.1) are calculated across all examples, regardless of the original dataset split. For the majority of scenarios, the SINGLE strategy outperforms the the MULTI approach in eliciting correct answers. Therefore, the SINGLE approach demonstrates higher agreement and tighter accuracy bounds, while the MULTI approach introduces more diverse responses and hallucinations with negligible impact on modal accuracy, allowing our simulation to generate more useful labels regarding query quality compared with a SINGLE approach.
@@
+ Calibration 84.7 73.5 69.2 85.5 78.5 73.0 88.1 78.9 76.1 83.1 78.2 70.1
   + $\tau=0.326$ 84.8 73.6 69.4 83.5 80.0 75.6 75.0 71.8 70.5 94.2 90.4 81.6
Table 4: HalluciBot Binary Evaluation Statistics. We report the Accuracy, F1, Precision, and Recall for all data splits. Probability threshold $\tau$ is computed along the closed interval $[0,1]$ in increments of $0.001$ to maximize the validation F1 score for the final model. The best ablation per base model is underlined, while the overall best performing model is in bold.
3.2 Converting Monte Carlo Estimates To Labels
Empirical Estimate. The probability of hallucination for a query $q_{0}$, denoted as $p_{h}(q_{0})$, can be empirically estimated based on the output $a_{i}\in\mathcal{A}$ of our Multi-Agent Monte Carlo simulation. We define the indicator function $\mathbb{I}$ to measure the incorrectness of an output $a_{i}$ with respect to the ground truth $y$ for query $q_{0}$.
$p_{h}\bigl{(}q_{0}\bigr{)}\approx\frac{1}{n+1}\sum\mathbb{I}\big{[}a_{i}\not=y\big{]}$
Binary Hallucination & Consensus Labels. To assess the propensity to hallucinate, we simplify the problem by considering two response values: whether $\mathit{q_{0}}$ produces any hallucination or not. Thus, we define the binary values for the probability of any hallucination as $p_{b}(q_{0})$. Furthermore, we craft a secondary consensus label $p_{c}(q_{0})$ that is a proxy for the agreement of the query. It maps the set of unique answers to $1$ if there is any disagreement, otherwise we assign 00 for perfect agreement ($1$ unique answer). Therefore, we can train a 2 head output Consensus model to predict the hallucination probability $p_{b}(q_{0})$, and if the query will cause confusion or consensus $p_{c}(q_{0})$.
$p_{b}\bigl{(}q_{0}\bigr{)}$ =$=\begin{cases}1&\text{if }p_{h}\bigl{(}q_{0}\bigr{)}>0\\ 0&\text{if }p_{h}\bigl{(}q_{0}\bigr{)}=0\\ \end{cases}$
$p_{c}\bigl{(}q_{0}\bigr{)}$ =$=\begin{cases}1&\text{if }\bigl{|}\{a_{i}\,|\,a_{i}\in\mathcal{A}\}\bigr{|}>1\\ 0&\text{if }\bigl{|}\{a_{i}\,|\,a_{i}\in\mathcal{A}\}\bigr{|}=1\\ \end{cases}$
3.3 How To Train a Classifier?
Once the Monte Carlo simulation is complete for our training corpus composed of 369,837 queries spanning 13 different datasets (Appendix C, Tables 1 & 13), we start training our classifier. These queries encompass Extractive, Multiple Choice, and Abstractive scenarios. Each scenario, with or without additional context, affects the hallucination rate of gpt-3.5-turbo. These simulated estimates are directly proportional to the approximated rates of hallucination $p_{h}$.
@@ 3.3 How To Train a Classifier?
    With a synthetic labeled set of queries $q_{0}$ and their rate of hallucinations $p_{h}(q_{0})$, we train an encoder-style RoBERTa (Liu et al. 2019) classifier to estimate the hallucination probability density from our Monte Carlo simulation.
[…]
    We ablate two versions: a binary model to estimate the propensity a query can hallucinate, and a consensus-aware model to also predict the expected agreement of outputs if sampled $n+1$ times.
Our experiments constrain the number of perturbations to $n=5$, and when including the original query and output, we can model the hallucination rate for $n+1=6$ modes. This translates to increments of $16.\overline{6}\%$ in hallucination rates.
Figure 5: Top: HalluciBot calibration curves with Brier Scores (BS), alongside the histogram of predicted probabilities. Bottom: Predicted hallucination labels juxtaposed against observed hallucination rates during our Monte Carlo simulation, with calibrated matrix below. We highlight $1$-$6$ as corresponding to the binary label “Yes - Hallucinatory” ($y=1$) during training. Notably, there is significant confusion in queries that are borderline ($1$, $2$) rather than majority hallucinatory prone ($3$-$6$).
How To Encode a Query’s Scenario? We conduct an ablation study to explore if incorporating the query’s scenario mitigates hallucinations. To create the prompt, we prepend the original query $q_{0}$ with either [EXTRACTIVE], [MULTIPLE CHOICE], or [ABSTRACTIVE], using the format <<{tag} $\mathtt{q_{0}}$}>>. Our hypothesis is based on recent research that highlights the use of RAG (Yan et al. 2024; Lewis et al. 2020; Guu et al. 2020) to alleviate hallucinations. The additional context provides valuable signals related to the hallucination rate of the original query. Furthermore, we apply this technique to distinguish our experimental results from reused datasets in different scenarios, such as SciQ (Johannes Welbl 2017) and SQuADv2 (Rajpurkar et al. 2016; Rajpurkar, Jia, and Liang 2018).
H4R: Downstream Modes For HalluciBot. Without HalluciBot’s feedback, typical query rewriting models have to act as both an implicit critic and a generator. As a proxy reward model, HalluciBot’s probabilistic feedback on a query’s quality, given the dual prediction heads for hallucination and consensus, can guide a query rewriting process using an independent gpt-3.5-turbo LLM, before output generation. In essence, HalluciBot provides the following downstream modes for handling potentially hallucinatory queries:
  1. 1.
    Rewrite Mode: A single-shot iterative rewrite of queries classified as hallucinatory.
  2. 2.
    Rank Mode: Generating $N$ intermediate perturbations, sorted by HalluciBot’s class probabilities for fine-grained scoring. In our implementation, the number of outputs were controlled by the number of chat completion choices in gpt-3.5-turbo’s API call.
  3. 3.
    Route Mode: For Abstractive or Extractive queries classified as hallucinatory, testing if switching the scenario (e.g. between RAG or direct inference) generates more robust classifications and generations.
The query rewriting prompt is found in Appendix Listing 1.
[…]
(B) [HB] (C) [HB]
(A) Naive Rewrite Informed Single Rewrite Best-of-N Rewrite
$+$ Class Transitions 6.5 $+$ Class Transitions 30.2 $+$ Class Transitions 50.6
-- Class Transitions 3.2 Rewrite Accuracy Rewrite Accuracy
Unneeded Rewrites 46.6 Top-5 94.3 Top-5 95.2
Similarity Score 46.9 Similarity Score 47.4
(D) Assuming HB Ratiocinate (E) [HB w/ Consensus] (F) [HB w/ Consensus]
Naive Rewrite (Baseline) Informed Single Rewrite Best-of-N Rewrite
$+$ Class Transitions 14.8 $+$ Class Transitions 31.9 $+$ Class Transitions 51.4
Rewrite Accuracy Rewrite Accuracy Rewrite Accuracy
Top-5 92.9 Top-5 90.2 Top-5 95.7
Similarity Score 41.7 Similarity Score 57.5 Similarity Score 55.9
Table 5: Query generation metrics under each HalluciBot (HB) strategy. Multiple Choice queries were evaluated on a soft accuracy criterion where the score is +1 if any of the $n$ generations match the ground truth. For Abstractive queries the average cosine similarity score between the ground truth and the $n$ generation outputs is reported. The embedding vectors for similarity computation are obtained using all-MiniLM-L6-v2 (Wang et al. 2020; Reimers and Gurevych 2019).
Figure 6: Class probability of queries that were rewritten and reclassified to be non-hallucinatory.
4 Experimental Setup
Dataset Coverage & Scenario Split. Our experiments include 13 datasets (Table 1) divided into 3 scenarios: Extractive, Multiple Choice, and Abstractive. To evaluate the impact of context, we use SQuADv2 (Rajpurkar et al. 2016; Rajpurkar, Jia, and Liang 2018) to simulate RAG (Lewis et al. 2020; Guu et al. 2020). To assess the effect of multiple choice queries, we repurposed TruthfulQA (Lin, Hilton, and Evans 2022) and SciQ (Johannes Welbl 2017) for two experiments: one where the output agents select from the choices or context, and another where LLM agents generate outputs without context. We maintain the original train, validation, and test splits across scenarios to prevent information leakage to HalluciBot. Prompt templates for each gpt-3.5-turbo agent can be found in App. Table 8. All LLM agents share the same set of parameters, as described in App. Table 6.
HalluciBot Training Parameters & Environment. We employed HuggingFace’s Trainer class with the Adam optimizer (Kingma and Ba 2017) for training, reporting efficiency and training times in App. Table 11. All experiments were conducted on an AWS EC2 instance with a single GPU (App. Table 7). HalluciBot is fine-tuned from both pretrained BERT (Devlin et al. 2018) and RoBERTa (Liu et al. 2019) models (App. Table 12). To address label imbalance, we employed a weighted class loss where each class weight is assigned to its inverted frequency in the training set. The train, validation, and test splits follow the original divisions of the datasets. Specifically, there are 302,492 training, 44,491 validation, and 22,854 testing samples. The distribution of labels across these splits is summarized in Table 2, and fine-grained splits per set are enumerated in App. Table 18. We also apply Platt calibration (Vasilev and D’yakonov 2023; Guo et al. 2017; Niculescu-Mizil and Caruana 2005; Platt 1999) based on the validation logits to help ensure that the raw probabilities align better with the true class labels (Figure 5).
@@ 5 Analysis & Discussion
Ablation: Perturbations Induce Output Diversity. We examine the impact of perturbations on the robustness of gpt-3.5-turbo in question-answering tasks by comparing two strategies: Single Query, Multiple Outputs (SINGLE) and Single Query, Multiple Perturbations, Single Output (MULTI). In the SINGLE strategy, we sample $n+1$ outputs from the original query $q_{0}$. In the MULTI strategy, $n$ perturbations of the original query $q_{0}$ are used, and each perturbation $q_{i}$ is answered once. Table 3 shows that while baseline accuracy remains consistent, the lower bound accuracy drops by $12.1$ points in the MULTI setting. Additionally, agreement metrics, as indicated by Fleiss’s $\kappa$, decrease by $12.5$ points, indicating reduced consistency. In summary, (1) the SINGLE strategy results in higher agreement and lower-bound accuracy while (2) the MULTI strategy increases response diversity and hallucination rates but offers a slight improvement in upper-bound accuracy for Extractive and Multiple Choice scenarios. This suggests that perturbations can enhance query quality by introducing necessary diversity, despite minor variations in modal accuracy.
Ratiocinate: Can HalluciBot Detect Hallucinatory Queries? Differentiating the scenario in HalluciBot’s prompt yielded a strong $+10.3\%$ increase in validation F1 score. The calibrated, threshold-tuned RoBERTa-base HalluciBot in Table 4 achieves a test accuracy of $69.5\%$ with a macro F1-score of $76.0\%$. Further breaking down the results in Figure 5, calibrating our models with Platt scaling improves the discriminating power for borderline queries, where the observed number of hallucinations was minimal ($y\in\{1,2\}$). Finally, HalluciBot demonstrates strong recall scores ($89.0\%$ validation, $82.6\%$ testing) to effectively flag risky queries that are likely to generate at least one hallucination during inference. The importance of HalluciBot as a ratiocinating process can be seen in [Table 5 (A)] under a naive rewriting strategy. Without HalluciBot, a naive rewrite strategy has the potential to convert queries originally estimated to be non-hallucinatory to hallucinatory (negative class transition), because there is no mechanism to differentiate queries. With HalluciBot restricting the test set to only potentially hallucinatory queries (11.2K samples), a naive rewrite [Table 5 (D)] can only enact positive class transitions ($+14.8\%$), converting queries originally estimated to hallucinate to non-hallucinatory. Furthermore, HalluciBot acting as an arbitrator can prevent computationally expensive rewrite calls for $46.6\%$ of the test set (10.2K samples deemed to be non-hallucinatory).
Rewrite: As a Feedback Mechanism. HalluciBot’s feedback allows us to generate a more informed query [Table 5 (B)] resulting in better class transition probabilities than an uninformed rewrite strategy [Table 5 (D)]. This translates to a $14.8\%$ positive class transition and a $1.4\%$ increase in Multiple Choice accuracy as well as $5.2\%$ improvement in generation similarity for Abstractive queries. Utilizing consensus information during the query rewriting process [Table 5 (E)] generates a slightly larger positive class transition ($31.9\%$ vs. $30.2\%$) than without [Table 5 (B)].
Rank: Best-of-N Rewrite. A Best-of-N rewrite strategy [Table 5 (C, F)] demonstrates a $19.5\%$ and $20.4\%$ gain in positive class transitions in both experiments over a single rewrite [Table 5 (B, E)]. Therefore, HalluciBot’s estimated probabilities can be used as a proxy reward model when ranking $n$ sample perturbations. The violin plots in Figure 6 shows that HalluciBot is able to select better queries in the Best-of-N rewrite setting than with a single rewrite, with a higher median predicted non-hallucinatory probability ($79.5\%$ (C) vs. $78.4\%$ (B); $80.4\%$ (F) vs. $76.5\%$ (E)). This means that HalluciBot evaluated the rewritten queries to be non-hallucinatory with greater probability. All rewrites are single-shot without subsequent tuning or iterations.
Route: Abstractive to Extractive. The test set had 948 Abstractive queries that were classified to be hallucinatory. Conditioned on this information, switching the scenario to Extractive results in a $+60.0\%$ positive class transition. In contrast, a rewriting process without a scenario change has a much smaller $9.7\%$ class transition. With HalluciBot’s ability to distinguish scenarios, it can help determine whether direct inference or RAG is more effective for a particular query.